ROTATIONAL SPECTRUM OF CYANOPHOSPHINE, H_{2}PCN

\underline{LU KANG}, STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459; M. C. McCARTHY, and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138; and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138.

The a type transitions of the microwave rotational spectra of cyanophosphine, H_{2}PCN, have been investigated in selected frequency regions between 10 and 42.5 GHz by Fourier Transformation Microwave (FTMW) spectroscopy. Rotational, centrifugal distortion and ^{14}N nuclear quadrupole coupling constants as well as the spin rotation coupling constants of phosphorus have been determined. Density functional theory level ab initio calculations were performed to predict the molecular constants, and the predicted values are in good agreement with our experimentally determined results. The 13C and ^{15}N isotopomer transitions were also observed. The derived r_{0} structure is quite comparable to the calculated H_{2}PCN equilibrium geometry.